# Mutual Influence of Symmetries and Topological Field Theories

### v1

@@ 1. Introduction
Another place where higher categories arise is within the framework of our modern understanding of symmetries for quantum field theories. It is believed that the collection of symmetry operators and their composition rules assemble into some higher category. This paper aims to expand on this conventional wisdom by considering what can happen if we allow ourselves the freedom of stacking a quantum field theory by topological quantum field theories. We will give an example where the TQFT can interact with the categorical symmetry in an interesting manner, changing the structure of the symmetry, but leaving the QFT invariant. [^1] As we will elaborate in a later section, the QFT is technically subtly modified, but the modifications cannot be detected by local observables. As we move into higher dimensions and gain access to nontrivial TQFTs, it becomes natural to ask whether we should broaden our notion of “equivalence for a theory" to go beyond the traditional scope of automorphisms and stacking by invertible TQFTs.
@@
(Theorem 2.9) We show that the periodicity of $\varpi$ depends on the properties of $G_{b}$ and $\kappa$.
  • if $\kappa=0$: $\varpi$ experiences no shift under the stacking and condensing procedure
  • if $\kappa\neq 0$ and $\mathrm{Sq}^{1}\kappa=0$: $\varpi$ experiences a 2-periodic shift. [^3] In particular a single iteration sends $(\alpha,\beta,\gamma)\mapsto(\alpha+\kappa,\,\beta+\kappa\cup_{1}\alpha,\,\gamma+G_{\kappa}(\alpha))$ where $G_{\kappa}(\alpha)$ is some $\mathds{C}/\mathds{Z}$ valued function such that $G_{\kappa}(\alpha)+G_{\kappa}(\alpha+\kappa)=0$.
@@ Acknowledgments
It is a pleasure to thank Theo Johnson-Freyd, Ryohei Kobayashi, Ho Tat Lam, Dmitri Nikshych, Abhinav Prem, Xiao-Gang Wen, and Yunqin Zheng for helpful conversations. This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). Part of this work was completed at the KITP and benefited from interactions with the participants of the program “Generalized Symmetries in Quantum Field Theory: High Energy Physics, Condensed Matter, and Quantum Gravity”. The author M.Y. is supported by the EPSRC Open Fellowship EP/X01276X/1. The author D.T. is supported by the Simons Collaboration for Global Categorical Symmetries and the Nova Scotia Graduate Scholarship program.
@@ 2.1. Preliminaries
While our discussion has centered on twisted spin structures arising from vector bundle, this is not strictly necessary. In general, for $X$ a space, with $a\in H^{1}(X;\mathds{Z}/2)$, and $b\in H^{2}(X;\mathds{Z}/2)$. The $(X,a,b)$-twisted spin structure on a vector bundle $E\to M$ is data of a map $f\colon M\to X$ and trivializations of $w_{1}(E)+f^{*}(s)$ and $w_{2}(E)+w_{1}(E)^{2}+f^{*}(\omega)$, but we do not assume there is a vector bundle $V\rightarrow X$ such that $s=w_{1}(V)$ and $b=w_{2}(V)$. This type of twisted spin structure which just considers cohomology classes was coined the “fake vector bundle” twist in [DY23a]. It is still possible even in the fake vector bundle case to define Thom spectra, following [ABG+14a, ABG+14b].
@@
[…]
@@
Importantly, the groupoid $\mathbf{2sVect}^{\times}$ furthermore has a non-trivial automorphism given by fermion parity, i.e. $B\mathds{Z}/2\simeq\mathcal{A}ut^{br}(\mathbf{sVect})$, which acts by $-1$ on the odd vector spaces. The action of $B\mathds{Z}/2$ on $\mathbf{2sVect}^{\times}$ is given by the following fiber sequence
[…]
@@ A $n$-cocycle in twisted supercohomology is the data of a triple $(\alpha,\beta,\gamma)$ where:
(11) $\pi_{0}$&$\pi_{1}$&$\pi_{2}$&$\pi_{3}$\\ \hline\cr\\[-13.00005pt] 1&$\mathds{Z}/2\oplus\mathds{Z}/2$&$\mathds{Z}/2$&$\mathds{C}^{\times}$\\ \hline\cr\end{tabular}\,.$
@@
[…]
@@ 2.2. Shifting the cocycle
    Objects $1$ $f$ $e$ $m$
    Spin 00 $\frac{1}{2}$ $\frac{n}{16}$ $\frac{n}{16}$
@@
    Objects $1$ $f$ $a$ $\overline{a}$
    Spin 00 $\frac{1}{2}$ $\frac{n}{16}$ $\frac{n}{16}$
@@
    Objects $1$ $f$ $\sigma$
    Spin 00 1 $\frac{n}{16}$
@@
where the multiplication between two cochains means applying $\cup$. Notice that $d\beta^{\prime}=d\beta+d\Delta\beta$, hence $\Delta\beta$ takes a form such that $d\Delta\beta=\kappa\alpha+\alpha\kappa$. In particular, $\Delta\beta=\kappa\cup_{1}\alpha$ [^7] The expression for $\Delta\beta$ can be found universally, which means that there are no contributions from other closed forms. Applying this procedure again would amount to letting $\beta^{\prime\prime}=\beta+\Delta\beta+\Delta\beta^{\prime}$, and finding the form of $\Delta\beta^{\prime}$. In particular, $\Delta\beta^{\prime}=\kappa\cup_{1}(\alpha+\kappa)$, and therefore $\beta^{\prime\prime}=\beta+\kappa\cup_{1}\kappa=\beta+\mathrm{Sq}^{1}\kappa$. If $\mathrm{Sq}^{1}\kappa$ is trivial, then $\beta$ is 2-periodic, and if $\mathrm{Sq}^{1}\kappa$ is nontrivial, then $\beta$ is 4-periodic. The 4-periodicity appears essentially because iterating this process twice more will generate another copy of $\mathrm{Sq}^{1}\kappa$ that cancels the first and takes us back to $\beta$.
@@
Here, the initial state $|\psi\rangle_{d}$ is said to be equivalent to the final state $|\psi^{\prime}\rangle_{d}$ if it is possible to tensor $|\psi\rangle_{d}$ with some locally trivial resource $|\text{triv}\rangle_{d-1}$, and evolve under the operator $\mathcal{U}$ to the state $|\psi^{\prime}\rangle_{d}$, tensored also with a trivial resource. The resource state, or ancilla, $|\text{triv}\rangle_{d-1}$ is a state that is allowed to be inserted in a $d-1$-dimensional submanifold. The resource is local in the sense that it only occupies a submanifold rather than being allowed to extend to the entire $d$-manifold, and it is trivial in the sense that it can be connected to the trivial product state by some finite depth local unitary.
@@ 3. Interactions of the Symmetry and TQFT
Here, $\mathcal{W}itt$ is the Witt group of non-degenerate braided fusion categories [DMNO13], and $\mathcal{W}itt(\mathcal{E})$ is the group of non-degenerate braided fusion categories over $\mathcal{E}$ i.e. elements are a Witt class $[\mathcal{C}]$ for $\mathcal{C}$ a braided fusion category with $\mathcal{Z}_{(2)}(\mathcal{C})\cong\mathcal{E}$, and composition with respect to another class $[\mathcal{D}]$ is given by $[\mathcal{C}\boxtimes_{\mathcal{E}}\mathcal{D}]$ up to $\mathcal{E}$-Witt equivalence. The second map in Equation (24) is given by base change, sending $[\mathcal{A}]\mapsto[\mathcal{A}\boxtimes\mathcal{E}]$. In the case of $\mathcal{E}=\mathbf{sVect}$ we have a sequence
@@ 3.1. The (3+1)d TQFT for Fusion 2-Categories
[…]

### v2

@@ 1. Introduction
Another place where higher categories arise is within the framework of our modern understanding of symmetries for quantum field theories. It is believed that the collection of symmetry operators and their composition rules assemble into some higher category. This paper aims to expand on this conventional wisdom by considering what can happen if we allow ourselves the freedom of stacking a quantum field theory by topological quantum field theories. We will give an example where the TQFT can interact with the categorical symmetry in an interesting manner, changing the structure of the symmetry, but leaving the QFT invariant. [^1] As we will elaborate in a later section, the QFT is technically subtly modified, but the modifications cannot be detected by local observables. As we move into higher dimensions and gain access to nontrivial TQFTs, it becomes natural to ask whether we should broaden our notion of “equivalence for a theory” to go beyond the traditional scope of automorphisms and stacking by invertible TQFTs.
@@
(Theorem 2.9) The periodicity of $\varpi$ depends on the following properties of $G_{b}$ and $\kappa$.
  • if $\kappa=0$: $\varpi$ experiences no shift under the stacking and condensing procedure;
  • if $\kappa\neq 0$ and $\mathrm{Sq}^{1}\kappa=0$: $\varpi$ experiences a 2-periodic shift; [^3] In particular a single iteration sends $(\alpha,\beta,\gamma)\mapsto(\alpha+\kappa,\,\beta+\kappa\cup_{1}\alpha,\,\gamma+G_{\kappa}(\alpha))$ where $G_{\kappa}(\alpha)$ is some $\mathds{C}/\mathds{Z}$ valued function such that $G_{\kappa}(\alpha)+G_{\kappa}(\alpha+\kappa)=0$.
@@ Acknowledgments
It is a pleasure to thank Theo Johnson-Freyd, Ryohei Kobayashi, Ho Tat Lam, Dmitri Nikshych, Abhinav Prem, Xiao-Gang Wen, and Yunqin Zheng for helpful conversations. This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). Part of this work was completed at the KITP and benefited from interactions with the participants of the program “Generalized Symmetries in Quantum Field Theory: High Energy Physics, Condensed Matter, and Quantum Gravity”. The author M.Y. is supported by the EPSRC Open Fellowship EP/X01276X/1. The author D.T. is supported by the Simons Collaboration for Global Categorical Symmetries and the Nova Scotia Graduate Scholarship program. No authors have competing interests to declare that are relevant to the content of this article. There is no data available for this article to declare.
@@ 2.1. Preliminaries
While our discussion has centered on twisted spin structures arising from vector bundle, this is not strictly necessary. In general, for $X$ a space, with $a\in H^{1}(X;\mathds{Z}/2)$, and $b\in H^{2}(X;\mathds{Z}/2)$. The $(X,a,b)$-twisted spin structure on a vector bundle $E\to M$ is data of a map $f\colon M\to X$ and trivializations of $w_{1}(E)+f^{*}(s)$ and $w_{2}(E)+w_{1}(E)^{2}+f^{*}(\omega)$, but we do not assume there is a vector bundle $V\rightarrow X$ such that $s=w_{1}(V)$ and $b=w_{2}(V)$. This type of twisted spin structure which just considers cohomology classes was coined the “fake vector bundle” twist in [DY23a] see also [DY24] for an application. It is still possible even in the fake vector bundle case to define Thom spectra, following [ABG+14a, ABG+14b].
@@
The way of realizing supercohomology that was just explained can be seen as the dual of superhomology which we denote $\tau_{\leq 2}\mathit{ko}$, which can be viewed as a truncation of $\mathit{ko}$ by the following short exact sequences:
(7) $\tau_{\geq 3}\mathit{ko}\longrightarrow\mathit{ko}\longrightarrow\tau_{\leq 2}\mathit{ko}\,.$
where $\tau_{\geq k}$ denotes the homotopy groups in degree $k$ and above. Using this short exact sequence it is possible to develop a modified Adams spectral sequence to conduct computations for twisted $\tau_{\leq 2}\mathit{ko}$-homology. Examples of computations with applications to anomalies, will appear in [DYYa, DYYb].
@@
Importantly, the groupoid $\mathbf{2sVect}^{\times}$ furthermore has a non-trivial automorphism given by fermion parity, i.e. $B\mathds{Z}/2\simeq\mathcal{A}ut^{br}(\mathbf{sVect})$, which acts by 1-1 on the odd vector spaces. The action of $B\mathds{Z}/2$ on $\mathbf{2sVect}^{\times}$ is given by the following fiber sequence
(9) $B^{n-2}\mathbf{2sVect}^{\times}$$B^{n-2}\mathbf{2sVect}^{\times}/\!/(B\mathds{Z}/2)$$B^{2}\mathds{Z}/2\,.$}
@@ A $n$-cocycle in twisted supercohomology is the data of a triple $(\alpha,\beta,\gamma)$ where:
(11) $\begin{tabular}[]{|c|c|c|c|c|}\hline\cr$\pi_{0}$&$\pi_{1}$&$\pi_{2}$&$\pi_{3}$\\ \hline\cr\\[-13.00005pt] 1&$\mathds{Z}/2\oplus\mathds{Z}/2$&$\mathds{Z}/2$$\mathds{C}^{\times}$\\ \hline\cr\end{tabular}\,$.
@@
(12) $BG_{b}$$B^{2}\mathcal{Z}(\mathbf{2sVect})^{\times}$$B^{2}\mathds{Z}/2\,$$\scriptstyle{\kappa}$
@@ 2.2. Shifting the cocycle
    Objects 11 ff $e$ $m$
    Spin 0 $\frac{1}{2}$ $\frac{n}{16}$ $\frac{n}{16}$
@@
    Objects 11 ff $a$ $\overline{a}$
    Spin 0 $\frac{1}{2}$ $\frac{n}{16}$ $\frac{n}{16}$
@@
    Objects 11 ff $\sigma$
    Spin 0 1 $\frac{n}{16}$
@@
where the multiplication between two cochains means applying $\cup$. Notice that $d\beta^{\prime}=d\beta+d\Delta\beta$, hence $\Delta\beta$ takes a form such that $d\Delta\beta=\kappa\alpha+\alpha\kappa$. In particular, $\Delta\beta=\kappa\cup_{1}\alpha$. [^7] The expression for $\Delta\beta$ can be found universally, which means that there are no contributions from other closed forms. Applying this procedure again would amount to letting $\beta^{\prime\prime}=\beta+\Delta\beta+\Delta\beta^{\prime}$, and finding the form of $\Delta\beta^{\prime}$. In particular, $\Delta\beta^{\prime}=\kappa\cup_{1}(\alpha+\kappa)$, and therefore $\beta^{\prime\prime}=\beta+\kappa\cup_{1}\kappa=\beta+\mathrm{Sq}^{1}\kappa$. If $\mathrm{Sq}^{1}\kappa$ is trivial, then $\beta$ is 2-periodic, and if $\mathrm{Sq}^{1}\kappa$ is nontrivial, then $\beta$ is 4-periodic. The 4-periodicity appears essentially because iterating this process twice more will generate another copy of $\mathrm{Sq}^{1}\kappa$ that cancels the first and takes us back to $\beta$.
@@
Here, the initial state $|\psi\rangle_{d}$ is said to be equivalent to the final state $|\psi^{\prime}\rangle_{d}$ if it is possible to tensor $|\psi\rangle_{d}$ with some locally trivial resource $|\text{triv}\rangle_{d-1}$, and evolve under the operator $\mathcal{U}$ to the state $|\psi^{\prime}\rangle_{d}$, tensored also with a trivial resource. The resource state, or ancilla, $|\text{triv}\rangle_{d-1}$ is a state that is allowed to be inserted in a d1d-1-dimensional submanifold. The resource is local in the sense that it only occupies a submanifold rather than being allowed to extend to the entire $d$-manifold, and it is trivial in the sense that it can be connected to the trivial product state by some finite depth local unitary.
@@ 3. Interactions of the Symmetry and TQFT
Here, $\mathcal{W}itt$ is the Witt group of non-degenerate braided fusion categories [DMNO13], and $\mathcal{W}itt(\mathcal{E})$ is the group of non-degenerate braided fusion categories over $\mathcal{E}$ i.e. elements are a Witt class $\mathcal{C}]$ for $\mathcal{C}$ a braided fusion category with $\mathcal{Z}_{(2)}(\mathcal{C})\cong\mathcal{E}$, and composition with respect to another class $\mathcal{D}]$ is given by $\mathcal{C}\boxtimes_{\mathcal{E}}\mathcal{D}]$ up to $\mathcal{E}$-Witt equivalence. The second map in Equation (24) is given by base change, sending $\mathcal{A}]\mapsto[\mathcal{A}\boxtimes\mathcal{E}]$. In the case of $\mathcal{E}=\mathbf{sVect}$ we have a sequence
@@ 3.1. The (3+1)d TQFT for Fusion 2-Categories
(27) $\ldots$$\mathrm{rSH}^{*}(BG_{b})$$\mathrm{SH}^{*}(BG_{b})$$\mathrm{H}^{*-2}(BG_{b};\mathds{Z}/2)$$\ldots$$\scriptstyle{M}$